OpenStudy (anonymous):

what is the exact values of y(2) if given dy/dx sin(xy)

myininaya (myininaya):

that is dy/dx=sin(xy) ?

OpenStudy (anonymous):

yes

myininaya (myininaya):

integrate both sides with respect to x

myininaya (myininaya):

wait don't listen to me let me thing about this

OpenStudy (anonymous):

i think u are right but im have trouble integrating it

myininaya (myininaya):

i was gonna say treat y like a constant

OpenStudy (anonymous):

i did but i didnt get the right answer

OpenStudy (anonymous):

the final answer is y(2) 1.632.

myininaya (myininaya):

here is what i was going to do so we intergate dy/dx with respect to x and get y we integrate the other side with respect to x and get 1/y*(-cos(xy))

myininaya (myininaya):

+C

OpenStudy (anonymous):

how did you get 1/y

myininaya (myininaya):

find \[\int\limits_{}^{}\sin(3x)dx\]

OpenStudy (anonymous):

it would be -1/3cos3x

myininaya (myininaya):

right so if we are treating y like a constant then what is \[\int\limits_{}^{}\sin(xy)dx\]

OpenStudy (anonymous):

ok so what do i plug in y if given x=2

myininaya (myininaya):

but see i don't think this right i think we need to separtation of varaibles? but i don't know how to expand sin(xy) i'm thinking on this part

OpenStudy (anonymous):

this is about using eulers method

OpenStudy (anonymous):

by the way

myininaya (myininaya):

ok eulers method let me look into it chag do you have any thoughts?

myininaya (myininaya):

so we don't know what y(2) is right?

OpenStudy (anonymous):

right we want to find the actual value

myininaya (myininaya):

i'm learning right now about eulers method

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i know i have to find the integral bu i do not know what to plug in for y

OpenStudy (anonymous):

they gave me initial value but i do not think that would elp

OpenStudy (anonymous):

*help

myininaya (myininaya):

nath this is harder than what i would thought it would be

myininaya (myininaya):

http://www.math.tamu.edu/REU/comp/matode.pdf scroll to page 16 this might help